OpenStudy (anonymous):

find zeros of f(x) = 3x - 2??? HELP!

OpenStudy (anonymous):

@RoseDryer

OpenStudy (anonymous):

Well you set it to equal 0. 3x-2=0 3x=2 x=2/3 <---- That is your zero. So the point is \[(\frac{ 2 }{ 3 },0)\]

OpenStudy (anonymous):

So 2/3,0 is the answer, correct? @RoseDryer

OpenStudy (anonymous):

If your looking for a point on a graph than yes.